1. The percent increase in profits, for the last five years at Cure Medico are: 5.61, 7.85, 11.22, 19.77 and 23.87.

a) Find the geometric percent increase.

b) Find the arithmetic percent increase.

2. The heights of a group of newly arrived plants in a nursery are given below (in centimeters): 10 15 14 16 25 13 17 23.
a) Compute the mean deviation of the heights.

3. A large group of management trainees was given two types of aptitude tests a quantitative aptitude test and a management aptitude test. The arithmetic mean score on the quantitative aptitude test was 400, with a standard deviation of 20. The mean was 50 and the standard deviation for the management aptitude test was 10. Compare the relative dispersion in the two groups.

4. A sample of ten customer care executive employed in the customer service department of a large telecommunication company received the following number of calls yesterday:
85, 78, 90, 97, 86, 72, 95, 89, 90, and 75.

a) Compute the coefficient of skewness using Pearson’s method.
b) What is your conclusion regarding the skewness of the data?
Answer #1

Answer:

1. The percent increase in profits, for the last five years at Cure Medico are: 5.61, 7.85, 11.22, 19.77 and 23.87.

a) Find the geometric percent increase.

11.84505

b) Find the arithmetic percent increase.

68.32/5 = 13.664

2. The heights of a group of newly arrived plants in a nursery are given below (in centimeters): 10 15 14 16 25 13 17 23.
a) Compute the mean deviation of the heights.

mean = 16.625

x

|x-mean|

10

6.625

15

1.625

14

2.625

16

0.625

25

8.375

13

3.625

17

0.375

23

6.375

Total

30.25

mean deviation = 30.25/8=3.78

3. A large group of management trainees was given two types of aptitude tests a quantitative aptitude test and a management aptitude test. The arithmetic mean score on the quantitative aptitude test was 400, with a standard deviation of 20. The mean was 50 and the standard deviation for the management aptitude test was 10. Compare the relative dispersion in the two groups.

Coefficient of variation for quantitative aptitude test = (20/400)*100 =5%

Coefficient of variation for management aptitude test = (10/50)*100 =20%

variation in management aptitude test is larger than variation in quantitative aptitude test.

4. A sample of ten customer care executive employed in the customer service department of a large telecommunication company received the following number of calls yesterday:
85, 78, 90, 97, 86, 72, 95, 89, 90, and 75.

a) Compute the coefficient of skewness using Pearson’s method.

n

10

mean

85.700

sample standard deviation

8.327

median

87.50

Skewness = 3(mean-median)/sd = 3*(85.7-87.5)/8.327

= -0.6485

b) What is your conclusion regarding the skewness of the data?

Since the Skewness < 0, the distribution is negatively skewed.